ZILL DIFFEQMODAP11 4.4.011. Solve the given differential equation by undetermined coefficients.
Answer #1

Given differential equation y – y + y = 8 teš The auxiliary equation is m? - m+ = 0 (n-1)=0 The complementary solution is y=